# Weitzenböck Remainder Spectrum on Rational Homogeneous Varieties

### 15 Mar 2026 v1

@@ 1 Introduction
gives rise to cross-terms involving the non-constant scalar curvature and the curvature of the connection $A$ on the auxiliary line bundle $\bf{L}$}.
@@
    If ${\mathcal{D}}_{A}\colon\Gamma^{\infty}({\mathcal{S}}(X_{P}))\to\Gamma^{\infty}({\mathcal{S}}(X_{P}))$ is the Dirac operator defined by the underlying Yang-Mills connection $A$ on some ${\bf{L}}\in{\rm{Spin}}^{c}(X_{P})$, then
@@
Let $(X_{P},\omega)$ be a rational homogeneous variety, such that $\omega$ is a $G$-invariant Kähler metric, and let ${\bf{L}}\in{\rm{Spin}}^{c}(X_{P})$ be a ${\rm{Spin}}^{c}$ structure with associated complex spinor bundle $\mathcal{S}(X_{P})$. If ${\mathcal{D}}_{A}\colon\Gamma^{\infty}({\mathcal{S}}(X_{P}))\to\Gamma^{\infty}({\mathcal{S}}(X_{P}))$ is the Dirac operator defined by the underlying Yang-Mills connection $A$ on $\bf{L}$}, then any eigenvalue $\lambda$ of ${\mathcal{D}}_{A}$ satisfies the following
@@ 2.2 Spinc Dirac operator and Weitzenböck remainder
There is a local description of this connection. Consider a local section $\tilde{u}\in\Gamma(P_{\operatorname{Spin}^{c}(n)}(M))$ such that $\Lambda(\tilde{u})=(u,\gamma)$, where $u=(e_{1},\dots,e_{n})$ is a local orthonormal frame of $TM$ and $\gamma\in\Gamma(L)$ is a local section. A local spinor field can be expressed as $\Psi=[\tilde{u},\psi]$, where $\psi$ is a function with values in $\Delta_{n}$. Then the connection acts locally as:
@@
$\forall Y\in\Gamma(TM)$, $\cdot$ denotes Clifford multiplication, and $A$ is the local connection 1-form.
@@
In the above theorem we consider the curvature $F_{A}$ of $A$ acting via Clifford multiplication extended to differential forms [FRI00].
@@
As we see from the Schrödinger-Lichnerowicz formula, in order to study the spectrum of the ${\mathcal{D}_{A}}^{2}-{\bf{\Delta}}_{A}$, we need to deal with the operator $F_{A}\colon\Gamma^{\infty}(\mathcal{S}(M))\rightarrow\Gamma^{\infty}(\mathcal{S}(M))$. We shall explore this action in the particular case that $(M^{2m},g,J)$ is Kähler manifold. In this particular setting, choosing $A$ as Chern connection, it follows that $\theta=\frac{\sqrt{-1}F_{A}}{2\pi}$ is a $(1,1)$-form on $M$. For this particular class of $2$-forms on a Kähler manifold $(M^{2m},g,J)$, we have the following standard result.
@@ 2.3 Atiyah-Singer index theorem
where ${\widehat{A}}(M)$ is the $\widehat{A}-$class of $TM$ and $c_{1}(\bf{L})$ is the first Chern class of $\bf{L}$}.
@@ 4.2 Proof of Theorem B
(B) If ${\mathcal{D}}_{A}\colon{\mathcal{S}}(X_{P})\to{\mathcal{S}}(X_{P})$ is the Dirac operator defined by the underlying Yang-Mills connection $A$ on some ${\bf{L}}\in{\rm{Spin}}^{c}(X_{P})$, it follows that
@@
where $S(\omega)$ is the scalar curvature of the Riemannian metric underlying $\omega$ and $F_{A}={\rm{d}}A$ is the curvature of the Yang-Mills connection $A$. Now we observe that
@@
see for instance Eq. (3.29) and Eq. (3.2). Moreover, since $A$ is Yang-Mills, it follows that $\theta=\frac{\sqrt{-1}F_{A}}{2\pi}$ is a closed $G$-invariant real $(1,1)$-form. Therefore, it follows that all the eigenvalues of the operator ${\mathcal{D}}_{A}^{2}-{\bf{\Delta}}_{A}$ are constant. From Theorem A, we have

### 18 Mar 2026 v2

@@ 1 Introduction
gives rise to cross-terms involving the non-constant scalar curvature and the curvature of the connection AA on the auxiliary line bundle $\bf{L}$}.
@@
    If ${\mathcal{D}}_{A}\colon\Gamma^{\infty}({\mathcal{S}}(X_{P}))\to\Gamma^{\infty}({\mathcal{S}}(X_{P}))$ is the Dirac operator defined by the underlying Yang-Mills connection AA on some ${\bf{L}}\in{\rm{Spin}}^{c}(X_{P})$, then
@@
Let $(X_{P},\omega)$ be a rational homogeneous variety, such that $\omega$ is a $G$-invariant Kähler metric, and let ${\bf{L}}\in{\rm{Spin}}^{c}(X_{P})$ be a ${\rm{Spin}}^{c}$ structure with associated complex spinor bundle $\mathcal{S}(X_{P})$. If ${\mathcal{D}}_{A}\colon\Gamma^{\infty}({\mathcal{S}}(X_{P}))\to\Gamma^{\infty}({\mathcal{S}}(X_{P}))$ is the Dirac operator defined by the underlying Yang-Mills connection AA on $\bf{L}$}, then any eigenvalue $\lambda$ of ${\mathcal{D}}_{A}$ satisfies the following
@@ 2.2 Spinc Dirac operator and Weitzenböck remainder
There is a local description of this connection. Consider a local section $\tilde{u}\in\Gamma(P_{\operatorname{Spin}^{c}(n)}(M))$ such that $\Lambda(\tilde{u})=(u,\gamma)$, where $u=(e_{1},\dots,e_{n})$ is a local orthonormal frame of TMTM and $\gamma\in\Gamma(L)$ is a local section. A local spinor field can be expressed as $\Psi=[\tilde{u},\psi]$, where $\psi$ is a function with values in $\Delta_{n}$. Then the connection acts locally as:
@@
$\forall Y\in\Gamma(TM)$, $\cdot$ denotes Clifford multiplication, and AA is the local connection 1-form.
@@
In the above theorem we consider the curvature $F_{A}$ of AA acting via Clifford multiplication extended to differential forms [FRI00].
@@
As we see from the Schrödinger-Lichnerowicz formula, in order to study the spectrum of the ${\mathcal{D}_{A}}^{2}-{\bf{\Delta}}_{A}$, we need to deal with the operator $F_{A}\colon\Gamma^{\infty}(\mathcal{S}(M))\rightarrow\Gamma^{\infty}(\mathcal{S}(M))$. We shall explore this action in the particular case that $(M^{2m},g,J)$ is Kähler manifold. In this particular setting, choosing AA as Chern connection, it follows that $\theta=\frac{\sqrt{-1}F_{A}}{2\pi}$ is a $(1,1)$-form on $M$. For this particular class of 22-forms on a Kähler manifold $(M^{2m},g,J)$, we have the following standard result.
@@ 2.3 Atiyah-Singer index theorem
where ${\widehat{A}}(M)$ is the $\widehat{A}-$class of TMTM and $c_{1}(\bf{L})$ is the first Chern class of $\bf{L}$}.
@@ 4.2 Proof of Theorem B
(B) If ${\mathcal{D}}_{A}\colon{\mathcal{S}}(X_{P})\to{\mathcal{S}}(X_{P})$ is the Dirac operator defined by the underlying Yang-Mills connection AA on some ${\bf{L}}\in{\rm{Spin}}^{c}(X_{P})$, it follows that
@@
where $S(\omega)$ is the scalar curvature of the Riemannian metric underlying $\omega$ and $F_{A}={\rm{d}}A$ is the curvature of the Yang-Mills connection AA. Now we observe that
@@
see for instance Eq. (3.29) and Eq. (3.2). Moreover, since AA is Yang-Mills, it follows that $\theta=\frac{\sqrt{-1}F_{A}}{2\pi}$ is a closed $G$-invariant real $(1,1)$-form. Therefore, it follows that all the eigenvalues of the operator ${\mathcal{D}}_{A}^{2}-{\bf{\Delta}}_{A}$ are constant. From Theorem A, we have